OpenStudy (anonymous):

the second angle of a triangle garden is four times as large as th first. the third angle is 45 less than the sum of the other two angles how do i put this in a equation

OpenStudy (radar):

Let X = First angle 4X= The second angle. Given 4X+X-45=The third angle. Given What we know: The sum of the angles in a triangle =180 degrees. X+4X+4X+X-45=180 degrees simplifying: 10X-45=180 degrees adding 45 to both sides result in:10X=225 dividing both sides by 10 result in x=22.5 degrees (first angle) 4x=90 degrees second angle 4x+x-45=67.5 degrees, the third angle